Write a system of two linear equations with infinitely many solutions for the following.

A quantity of 13% acid solution is being diluted with some 3% acid solution. The final solution must be 8% acid and contain 90 mL. How much of each solution must be added to accomplish this?

a.) Write a system of equations that describes the situation.
b.) Solve the system.

Answer by checkley71(8403)

.13X+.03(90-X)=.08*90
.13X+2.7-.03X=7.2
.10X=7.2-2.7
.10X=4.5
X=4.5/.1
X=45 ML OF 13% SOLUTION IS USED.
90-45=45 ML OF 3% IS ALSO USED.
PROOF
.13*45+.03*45=.08*90
5.85+1.35=7.2
7.2=7.2